OpenStudy (anonymous):

Ok, this is quite a big one.. I just need to know what formulas to use for certain parts of the question: Matty the moth is on a turntable spinning at (2 3/8) rotations in 8 seconds. Bob is 11cm away from the center. A)Find the angular displacement of Bob. B)Find the angular velocity in rads per second. C)Find the distance that Bob traveled in 8 seconds. D)Find the linear velocity of Bob in cm per second.

OpenStudy (anonymous):

I already have the answers, just need the formulas and/or some examples of what to do -- studying.

OpenStudy (moongazer):

\[\omega=\theta/t\] \[l=s/t\]

OpenStudy (moongazer):

w is the angular velocity. l is the linear velocity. s is the arc length you can also make another equation for l by using those formula \[s=\theta r\]

OpenStudy (moongazer):

so l is also \[l=\omega r\]

OpenStudy (anonymous):

Thanks for the help

OpenStudy (moongazer):

also do not forget to convert revolutions (rev) to radians . theta should be in radians you can convert it by multiplying to 2pi

OpenStudy (anonymous):

\[\theta=revs \times2\pi rads\] Correct?

OpenStudy (moongazer):

yes, do you know why it is correct ?

OpenStudy (anonymous):

1rev = 2πrads

OpenStudy (moongazer):

correct :)

OpenStudy (anonymous):

I guess i have been paying enough attention in trig haha, thanks again

OpenStudy (moongazer):

your'e welcome :)